# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 8 Challenges of AI and communication for 6G
On the other hand, these advancements introduce complexities. The challenges of AI for the network include the strain on bandwidth and storage due to complex models. Conversely, the challenges of the network for AI require robust infrastructure to support advanced functionalities. The reliance on AI raises concerns about sustainability, as increased energy demands impact the environmental footprint. The requirement for real-time AI necessitates ultra-low latency, demanding sophisticated computational infrastructure. Security and privacy are at risk from potential attacks on sensitive data. Lastly, human-AI interactive collaboration presents ethical challenges, necessitating transparent oversight to ensure trust and accountability.
This section focuses on the multifaceted challenges faced by AI and communication for 6G from the abovementioned perspectives.
8.1 Challenges of AI For Network
The practiced deployment of AI in complex, time-varying wireless environments also challenges AI. At the same time, improving the interpretability of AI deployment in wireless environments, promoting people to understand AI, trust AI, and use AI will also be important in introducing AI into wireless communication.
@@ 8.1 Challenges of AI For Network
The generalization problem is widely prevalent in integrating AI with communication systems [291, 292, 293, 294]. Network models with poor generalization capabilities cannot often extend their learned knowledge to other variable domains. Consequently, such models favor one or a few data types, resulting in suboptimal performance when reasoning on different data types. This issue becomes particularly critical when the model is exposed to diverse datasets on which it was not explicitly trained, leading to a significant decline in its effectiveness.
The application of AI in wireless communications imposes even higher demands on model generalization [291, 292, 293, 295, 296, 297, 298, 299]. In wireless environments, users are situated in varying contexts, with factors such as the relative position to BSs and user mobility playing significant roles. This variability means that pre-trained AI models may lose effectiveness when confronted with new data that deviate from their training sets. In the dynamic and complex wireless environments characteristic of 6G networks, models with poor generalization capabilities will lack the potential for deployment and widespread use. Therefore, enhancing the generalization ability of AI models within 6G networks is a critical issue requiring industry and academia dedicated attention. Addressing this challenge will ensure that AI models maintain high performance and adaptability across diverse and evolving wireless scenarios, thus facilitating their successful integration into future communication systems.
@@
Robustness, derived from the concept of cybernetics [300], plays a central role in AI as it measures whether an AI model can maintain its efficacy and reliability in the face of real-world complexity and uncertainty. A robust AI model can keep the stability of its output in the face of minor deviations from the data. This property is critical in determining whether the model suits real-world scenarios.
In modern wireless communication systems, AI technologies such as deep learning are playing an increasingly important role[300]. For example, in the field of semantic communication, deep neural networks are widely used for critical tasks such as semantic coding, decoding and understanding[253, 256, 301]. However, the fragility of deep learning models and the open nature of wireless channels make these AI-powered communication systems extremely vulnerable to physical layer adversarial attacks. Attackers can mislead the semantic interpretation on the receiving end by injecting carefully designed perturbation signals, causing catastrophic consequences in certain critical domains that rely on highly reliable communication (e.g., autonomous driving, smart healthcare, etc.). In this context, the robustness of AI algorithms in communication systems is particularly important[300, 253, 256, 301]. Ensuring the robustness of AI algorithms in communication systems has become a critical issue.
Interpretability of AI
Deep learning is essentially a black box due to its highly nonlinear nature. The model learns through neurons, convolution kernel and other ways. People cannot understand the meaning, importance and fluctuation range of parameters in the neural network through basic statistical assumptions like linear regression parameters, which leads to unexplainability. For example, we only focus on the encoder and decoder output if we use AI for CSI feedback. We do not have a concrete understanding of how AI performs feature extraction and recovery, leading to the randomness of model training and inference performance that is difficult to control and increased unreliability against harsh, extreme environments.
Interpretable AI is not only a technical issue but also a complex issue involving trust, fairness, regulation, and human relations. Improving the explainability of AI and helping human users understand the decision-making process, working principle and potential bias of AI to build trust in AI is necessary to promote the broad application of AI in communication and even in various fields. The millions or even billions of parameters in deep neural networks and their nonlinear activation function form a highly complex decision boundary. This complexity makes the model’s decision-making process difficult to understand intuitively, thus posing significant challenges for interpretability.
8.2 Challenges of Network For AI
The inherent characteristics of wireless communication will have an impact on the performance of RAN in supporting AI services, mainly in five aspects: the dynamics of wireless transmission, the heterogeneity of wireless networks, the discrete nature of wireless nodes, the large-scale complexity of network development, and the scarcity of computing resources in wireless networks. In this paper, we research the characteristics of these five aspects to explore their impact on AI’s performance and deployment realization in wireless access networks and propose potential solutions.
@@ 8.2 Challenges of Network For AI
Constructing an AI execution environment in wireless networks inherently entails significant dynamic variations. The interference and loss in the wireless connection between terminals and BSs are impacted by the surrounding urban environment and changes in the operational states of other BSs, subsequently affecting the uplink and downlink connection rates. At the same time, terminals exhibit high mobility, potentially moving from the center to the edge of a cell, switching to another cell, or even leaving the wireless coverage area. Additionally, distinct from dedicated computing resources in IT clouds, wireless networks carry traditional mobile connectivity tasks, with the mobile traffic load on individual network elements continually fluctuating. BSs alternate between busy and idle periods, and terminals enter an idle state when there are no uplink or downlink services. These variations contribute to the dynamic nature of mobile wireless networks, resulting in slow model training convergence and adversely affecting model inference performance. It can lead to the failure or loss of partial information, such as gradients or losses, during AI model computations, further slowing down convergence and, in severe cases, causing divergence and failure to converge.
Acknowledging the influence of these dynamic wireless transmission factors, the authors of [302] designed a mitigation strategy for unreliable communication nodes. This approach leverages dynamic wireless channel state predictions to estimate the training completion latency for each user equipment’s local model. It avoids unreliable wireless communication nodes, prioritizing nodes with stable and high-speed wireless connections for training participation. When unstable or disconnected device nodes occur during training, the strategy advocates for their exclusion to ensure continuous and efficient AI training. Simulation results demonstrate that avoiding unreliable nodes facilitates stable convergence during AI model training.
@@
In wireless communication networks, a myriad of heterogeneities exist, encompassing communication heterogeneity arising from the varying geographical locations of nodes and the time-varying nature of communication network links; computational heterogeneity due to differences in the number, manufacturing processes, and architectures of node chips; data heterogeneity in terms of the volume and distribution of data possessed by nodes, influenced by factors such as storage capacity and application scenarios; and model heterogeneity stemming from varying task objectives that lead to the training of diverse models. These heterogeneities pose challenges to achieving efficient distributed learning.
This diverse heterogeneity influences both AI performance and convergence. Data heterogeneity results in disparities among local gradients computed by AI models, causing inconsistency in the direction of aggregated gradients, which can lead to model divergence or even non-convergence. Differences in computational power and communication capabilities lead to variations in AI model computation and information transmission efficiency, giving rise to the straggler effect. Furthermore, in AI learning, when the number of user devices is sufficiently large or the local models are adequately complex, the consumption of wireless resources during transmission becomes significant.
@@
The discrete nature of 6G mobile networks, characterized by highly heterogeneous and dynamically changing environments, poses significant challenges to the application of AI. This discreteness manifests through the rapid and unpredictable variations in network topology, user mobility patterns, and service demands, leading to inconsistent and fragmented data streams. Such conditions impede the ability of AI models to maintain accuracy and reliability, as the models must continuously adapt to a multitude of scenarios with limited contextual information. The resulting variability can degrade the performance of AI algorithms, particularly those relying on stable and predictable data inputs, thereby affecting tasks such as predictive analytics, resource allocation, and anomaly detection.
These challenges necessitate a concerted effort from academia and industry to develop AI models with enhanced robustness and generalization capabilities. This involves advancing research in adaptive learning techniques, transfer learning, and FL for the academic community to handle better the sporadic and diverse data inherent in 6G networks. For the industry, there is a pressing need to design and implement resilient AI-driven network management systems capable of real-time adaptation to fluctuating conditions. Addressing these challenges satisfactorily will be crucial for leveraging the full potential of AI in optimizing the performance and efficiency of 6G mobile networks.
@@
The large scale and complexity of 6G networks significantly affect the deployment and performance of AI systems. The many connected devices, diverse applications, and heterogeneous network environments introduce unprecedented data volume, variety, and velocity levels. AI models must process and analyze this extensive data in real time to make effective decisions, which is a daunting task due to the high-dimensional and dynamic nature of the information. The large-scale deployment of AI in 6G networks also necessitates distributed and decentralized processing architectures, complicating data aggregation and synchronization. Furthermore, the intricate interplay between various network components, such as edge and core networks, requires AI models to handle multifaceted interactions and dependencies, leading to increased computational overhead and latency. Consequently, AI algorithms’ scalability and efficiency are tested, making it challenging to maintain high performance and reliability across the entire network.
Addressing these challenges demands innovative approaches from both academia and industry. From the perspective of academia, there is a need to advance research in scalable AI techniques, including distributed learning, edge computing, and hierarchical models, to efficiently handle the massive and complex data in 6G networks. Developing AI algorithms capable of real-time processing and decision-making with minimal latency is crucial. Additionally, robust frameworks for integrating AI with existing network infrastructures must be explored to ensure seamless operation. The industry should focus on implementing scalable AI-driven solutions that can dynamically adapt to the network’s complexity and scale. This involves designing intelligent network management systems that can autonomously optimize resources, enhance security, and improve user experiences. Collaborative efforts between academia and industry will be essential to develop and deploy AI technologies that can fully leverage the capabilities of 6G networks, ensuring robust performance and efficient operation in an increasingly complex and large-scale environment.
@@
In large-scale and dynamic edge federated networks, the localized model training and transmission may not always succeed due to constraints on mobile devices’ power and computational resources, bandwidth limitations, wireless channel impairments, etc., directly compromising the performance of FL regarding model accuracy and convergence speed. Consequently, evaluating the trade-off relationship between cognitive performance and multi-dimensional resource allocation becomes imperative upon introducing distributed learning.
The authors of [303] focus on FL-enabled edge intelligence networks by leveraging Poisson Point Processes to construct an analytical model. The work delves into three specific scenarios, each deriving mathematical relationships between FL performance, consumed computational resources, and communication resources by analyzing the impact of Signal-to-Interference-plus-Noise Ratio, SNR, and faulty transmissions on local and global models. This provides a theoretical foundation and basis for optimizing the framework design of intelligent access networks and allocating communication, computing power, and data resources.
@@ 8.3 Consideration of Sustainability
The imperative to reduce energy consumption in the context of AI and 6G integration is deeply rooted in the global consensus on carbon neutrality and peak carbon emissions. Through agreements like the Paris Agreement, the international community has committed to significant environmental goals, including halving carbon emissions by 2030. These commitments reflect an agreement on the urgent need to combat climate change and its potentially devastating impacts. AI’s role in enhancing 6G network performance is indisputable. The advanced capabilities of AI are crucial for managing the complex operations of 6G networks, which promise unprecedented speeds, low latency, and high reliability. However, the computational power required for AI comes at a high energy cost, which, if unchecked, could increase carbon emissions. This directly conflicts with the global movement towards reducing greenhouse gas emissions and achieving a sustainable, low-carbon future [304]. Therefore, the concept of ”green” development has become a core tenet of 6G evolution. The telecommunications industry is under increasing pressure to demonstrate environmental responsibility by minimizing its carbon footprint. This is a response to regulatory requirements and mandates and reflects consumer expectations and corporate social responsibility. As a result, striking a balance between performance and green objectives has emerged as one of the main challenges for future 6G development.
Techniques such as pruning, quantization, and using neural architecture search can help create leaner models that maintain high accuracy while reducing computational load [305]. Developing energy-aware AI models that dynamically adjust their energy consumption based on the current network demands can contribute to a more sustainable operation [306]. Another aspect is optimizing data processing and transmission within 6G networks [307]. Leveraging AI for intelligent traffic routing and predictive maintenance can reduce unnecessary energy expenditure. Furthermore, deploying AI at the network’s edge can help localize data processing, reducing the energy costs associated with data transfer across long distances. Integrating renewable energy sources into 6G infrastructure is also critical to the green 6G vision. The telecommunications industry can significantly reduce its reliance on fossil fuels and its overall carbon footprint by powering data centers and network operations with solar, wind, and other renewable energies.

### v3

@@ 7 Challenges of AI and communication for 6G
On the other hand, these advancements introduce complexities. The challenges of AI for the network include the strain on bandwidth and storage due to complex models. Conversely, the challenges of the NET4AI require robust infrastructure to support advanced functionalities. The reliance on AI raises concerns about sustainability, as increased energy demands impact the environmental footprint. The requirement for real-time AI necessitates ultra-low latency, demanding sophisticated computational infrastructure. Security and privacy are at risk from potential attacks on sensitive data. Lastly, human-AI interactive collaboration presents ethical challenges, necessitating transparent oversight to ensure trust and accountability. This section focuses on the multifaceted challenges faced by AI and communication for 6G from the abovementioned perspectives.
7.1 Challenges of AI4NET
Reliability of AI
The reliability of AI in networks poses a significant challenge due to the dynamic, heterogeneous, and often unpredictable nature of 6G communication systems [261]. AI models, which rely heavily on data-driven learning, can struggle to maintain consistent performance when exposed to conditions that deviate from their training environments. For instance, unexpected network anomalies, rare edge cases, or sudden shifts in user behavior patterns can lead to errors or even system failures [262]. Moreover, the non-deterministic nature of many AI algorithms makes it difficult to guarantee predictable outcomes, especially in time-sensitive or mission-critical applications. Factors such as data sparsity, incomplete information, and the presence of adversarial inputs further undermine the reliability of AI, as these systems may produce incorrect predictions or decisions under such conditions. Additionally, the distributed and decentralized nature of 6G networks introduces further complexity, as AI models must operate across diverse nodes with varying resource constraints and environmental conditions. This interconnectedness increases the risk of cascading failures, where an error in one part of the system can propagate and disrupt the entire network. These challenges highlight the fundamental difficulty in ensuring that AI systems remain reliable, robust, and consistent in the face of evolving and unpredictable network demands.
Stability of AI
The stability of AI in networks is a critical challenge, as it directly impacts the consistency and predictability of AI-driven decisions in dynamic and complex environments [263]. Stability refers to an AI system’s ability to produce reliable outputs when exposed to variations in input data, environmental conditions, or system parameters. In networked systems, this becomes particularly challenging due to the heterogeneity and unpredictability of real-world conditions [264]. For instance, fluctuating network traffic, sudden surges in user demand, or unexpected hardware failures can introduce variations that destabilize AI models. This instability can manifest as erratic predictions, oscillating decision patterns, or even complete system breakdowns. Furthermore, the non-linear and highly sensitive nature of many AI algorithms, such as DL models, makes them prone to overreacting to small changes in input data, leading to cascading errors. In distributed or decentralized networks, such instability can propagate across nodes, amplifying its impact and potentially disrupting the entire system. Additionally, the lack of transparency in many AI models (e.g., black-box models) complicates the diagnosis and mitigation of instability, as it is often unclear how or why a model’s output changes under varying conditions. These challenges highlight the inherent difficulty in ensuring stable and predictable AI performance in the face of the complex, evolving, and often adversarial nature of modern networks.
@@ 7.1 Challenges of AI4NET
The generalization problem is widely prevalent in integrating AI with communication systems [265, 266, 267, 268]. Network models with poor generalization capabilities cannot often extend their learned knowledge to other variable domains. Consequently, such models favor one or a few data types, resulting in suboptimal performance when reasoning on different data types. This issue becomes particularly critical when the model is exposed to diverse datasets on which it was not explicitly trained, leading to a significant decline in its effectiveness. The application of AI in wireless communications imposes even higher demands on model generalization [269, 270, 271, 272, 273]. In wireless environments, users are situated in varying contexts, with factors such as the relative position to BSs and user mobility playing significant roles. This variability means that pre-trained AI models may lose effectiveness when confronted with new data that deviate from their training sets. In the dynamic and complex wireless environments characteristic of 6G networks, models with poor generalization capabilities will lack the potential for deployment and widespread use. Therefore, enhancing the generalization ability of AI models within 6G networks is a critical issue requiring industry and academia dedicated attention. Addressing this challenge will ensure that AI models maintain high performance and adaptability across diverse and evolving wireless scenarios, thus facilitating their successful integration into future communication systems.
Interpretability of AI
DL is essentially a black box due to its highly nonlinear nature. The model learns through neurons, convolution kernel and other ways. People cannot understand the meaning, importance and fluctuation range of parameters in the neural network through basic statistical assumptions like linear regression parameters, which leads to unexplainability. For example, we only focus on the encoder and decoder output if we use AI for CSI feedback. We do not have a concrete understanding of how AI performs feature extraction and recovery, leading to the randomness of model training and inference performance that is difficult to control and increased unreliability against harsh, extreme environments. Interpretable AI is not only a technical issue but also a complex issue involving trust, fairness, regulation, and human relations. Improving the explainability of AI and helping human users understand the decision-making process, working principle and potential bias of AI to build trust in AI is necessary to promote the broad application of AI in communication and even in various fields. The millions or even billions of parameters in DNNs and their nonlinear activation function form a highly complex decision boundary. This complexity makes the model’s decision-making process difficult to understand intuitively, thus posing significant challenges for interpretability.
@@
Robustness, derived from the concept of cybernetics, plays a central role in AI as it measures whether an AI model can maintain its efficacy and reliability in the face of real-world complexity and uncertainty. A robust AI model can keep the stability of its output in the face of minor deviations from the data. This property is critical in determining whether the model suits real-world scenarios. In modern wireless communication systems, AI technologies such as DL are playing an increasingly important role[274]. For example, in the field of semantic communication, DNNs are widely used for critical tasks such as semantic coding, decoding and understanding[275]. However, the fragility of DL models and the open nature of wireless channels make these AI-powered communication systems extremely vulnerable to physical layer adversarial attacks. Attackers can mislead the semantic interpretation on the receiving end by injecting carefully designed perturbation signals, causing catastrophic consequences in certain critical domains that rely on highly reliable communication (e.g., autonomous driving, smart healthcare, etc.). In this context, the robustness of AI algorithms in communication systems is particularly important. Ensuring the robustness of AI algorithms in communication systems has become a critical issue.
7.2 Challenges of NET4AI
@@
Constructing an AI execution environment in wireless networks inherently entails significant dynamic variations. The interference and loss in the wireless connection between terminals and BSs are impacted by the surrounding urban environment and changes in the operational states of other BSs, subsequently affecting the uplink and downlink connection rates. At the same time, terminals exhibit high mobility, potentially moving from the center to the edge of a cell, hand off to another cell, or even leaving the wireless coverage area. Additionally, distinct from dedicated computing resources in IT clouds, wireless networks carry conventional mobile connectivity tasks, with the mobile traffic load on individual network elements continually fluctuating. BSs alternate between busy and idle periods, and terminals enter an idle state when there are no uplink or downlink services. These variations contribute to the dynamic nature of mobile wireless networks, resulting in slow model training convergence and adversely affecting model inference performance. It can lead to the failure or loss of partial information, such as gradients or losses, during AI model computations, further slowing down convergence and, in severe cases, causing divergence and failure to converge. Acknowledging the influence of these dynamic wireless transmission factors, the authors of [276] designed a mitigation strategy for unreliable communication nodes. This approach leverages dynamic wireless channel state predictions to estimate the training completion latency for each user equipment’s local model. It avoids unreliable wireless communication nodes, prioritizing nodes with stable and high-speed wireless connections for training participation. When unstable or disconnected device nodes occur during training, the strategy advocates for their exclusion to ensure continuous and efficient AI training. Simulation results demonstrate that avoiding unreliable nodes facilitates stable convergence during AI model training.
@@
In wireless communication networks, a myriad of heterogeneities exist, encompassing communication heterogeneity arising from the varying geographical locations of nodes and the time-varying nature of communication network links; computational heterogeneity due to differences in the number, manufacturing processes, and architectures of node chips; data heterogeneity in terms of the volume and distribution of data possessed by nodes, influenced by factors such as storage capacity and application scenarios; and model heterogeneity stemming from varying task objectives that lead to the training of diverse models. These heterogeneities pose challenges to achieving efficient distributed learning. This diverse heterogeneity influences both AI performance and convergence. Data heterogeneity results in disparities among local gradients computed by AI models, causing inconsistency in the direction of aggregated gradients, which can lead to model divergence or even non-convergence. Differences in computational power and communication capabilities lead to variations in AI model computation and information transmission efficiency, giving rise to the straggler effect. Furthermore, in AI learning, when the number of user devices is sufficiently large or the local models are adequately complex, the consumption of wireless resources during transmission becomes significant.
@@
The discrete nature of 6G mobile networks, characterized by highly heterogeneous and dynamically changing environments, poses significant challenges to the application of AI. This discreteness manifests through the rapid and unpredictable variations in network topology, user mobility patterns, and service demands, leading to inconsistent and fragmented data streams. Such conditions impede the ability of AI models to maintain accuracy and reliability, as the models must continuously adapt to a multitude of scenarios with limited contextual information. The resulting variability can degrade the performance of AI algorithms, particularly those relying on stable and predictable data inputs, thereby affecting tasks such as predictive analytics, resource allocation, and anomaly detection. These challenges necessitate a concerted effort from academia and industry to develop AI models with enhanced robustness and generalization capabilities. This involves advancing research in adaptive learning techniques, transfer learning, and FL for the academic community to handle better the sporadic and diverse data inherent in 6G networks. For the industry, there is a pressing need to design and implement resilient AI-driven network management systems capable of real-time adaptation to fluctuating conditions. Addressing these challenges satisfactorily will be crucial for leveraging the full potential of AI in optimizing the performance and efficiency of 6G mobile networks.
@@
The large scale and complexity of 6G networks significantly affect the deployment and performance of AI systems. The many connected devices, diverse applications, and heterogeneous network environments introduce unprecedented data volume, variety, and velocity levels. AI models must process and analyze this extensive data in real time to make effective decisions, which is a daunting task due to the high-dimensional and dynamic nature of the information. The large-scale deployment of AI in 6G networks also necessitates distributed and decentralized processing architectures, complicating data aggregation and synchronization. Furthermore, the intricate interplay between various network components, such as edge and CNs, requires AI models to handle multifaceted interactions and dependencies, leading to increased computational overhead and latency. Consequently, AI algorithms’ scalability and efficiency are tested, making it challenging to maintain high performance and reliability across the entire network. Addressing these challenges demands innovative approaches from both academia and industry. From the perspective of academia, there is a need to advance research in scalable AI techniques, including distributed learning, edge computing, and hierarchical models, to efficiently handle the massive and complex data in 6G networks. Developing AI algorithms capable of real-time processing and decision-making with minimal latency is crucial. Additionally, robust frameworks for integrating AI with existing network infrastructures must be explored to ensure seamless operation. The industry should focus on implementing scalable AI-driven solutions that can dynamically adapt to the network’s complexity and scale. This involves designing intelligent network management systems that can autonomously optimize resources, enhance security, and improve user experiences. Collaborative efforts between academia and industry will be essential to develop and deploy AI technologies that can fully leverage the capabilities of 6G networks, ensuring robust performance and efficient operation in an increasingly complex and large-scale environment.
@@
In large-scale and dynamic edge federated networks, the localized model training and transmission may not always succeed due to constraints on mobile devices’ power and computational resources, bandwidth limitations, wireless channel impairments, etc., directly compromising the performance of FL regarding model accuracy and convergence speed. Consequently, evaluating the trade-off relationship between cognitive performance and multi-dimensional resource allocation becomes imperative upon introducing distributed learning. The authors of [277] focus on FL-enabled edge intelligence networks by leveraging Poisson Point Processes to construct an analytical model. The work delves into three specific scenarios, each deriving mathematical relationships between FL performance, consumed computational resources, and communication resources by analyzing the impact of Signal-to-Interference-plus-Noise Ratio, SNR, and faulty transmissions on local and global models. This provides a theoretical foundation and basis for optimizing the framework design of intelligent access networks and allocating communication, computing power, and data resources.
@@ 7.3 Consideration of Sustainability
The imperative to reduce energy consumption in the context of AI and 6G integration is deeply rooted in the global consensus on carbon neutrality and peak carbon emissions [278]. Through agreements like the Paris Agreement, the international community has committed to significant environmental goals, including halving carbon emissions by 2030. The advanced capabilities of AI are crucial for managing the complex operations of 6G networks, which promise unprecedented speeds, low latency, and high reliability. However, the computational power required for AI comes at a high energy cost, which, if unchecked, could increase carbon emissions. This directly conflicts with the global movement towards reducing greenhouse gas emissions and achieving a sustainable, low-carbon future. Therefore, the concept of “Green” development has become a core tenet of 6G evolution.
Particularly, the energy consumption of AI is not only restricted to the communication network itself but also pertains to cloud computing centers and edge computing devices. For example, the computational requirements of LLM (such as GPT-4) and complex AI tasks (such as RL) have increased substantially [279]. In the training stage, it may consume several hundreds to thousands of MWh of energy, while the energy consumption in the inference stage is directly related to the invocation frequency, model scale, and real-time requirements. Meanwhile, the leasing costs of cloud computing resources are affected by computing power, usage duration, storage requirements, and market prices. Users typically bear these costs through pay-as-you-go, reserved instance, or subscription models and may indirectly pass them on to end consumers through subscription fees or service charges. Regarding the evaluation of energy efficiency, it can be analyzed from multiple aspects. The energy efficiency of the communication system itself is usually measured by indicators such as energy per bit or spectral efficiency. For the energy consumption assessment of cloud computing centers and the edge, a more comprehensive perspective is needed. For instance, the energy efficiency of a data center can be evaluated by the power usage effectiveness metric, which reflects the ratio of the total energy consumption of the data center to the energy consumption of IT equipment. The closer the power usage effectiveness is to 1, the higher the energy utilization efficiency.
To address these challenges, techniques such as pruning, quantization, and using neural architecture search can help create leaner models that maintain high accuracy while reducing computational load [280]. Developing energy-aware AI models that dynamically adjust their energy consumption based on the current network demands can contribute to a more sustainable operation [281]. Another aspect is optimizing data processing and transmission within 6G networks [282]. Leveraging AI for intelligent traffic routing and predictive maintenance can reduce unnecessary energy expenditure. Furthermore, deploying AI at the network’s edge can help localize data processing, reducing the energy costs associated with data transfer across long distances. Integrating renewable energy sources into 6G infrastructure is also critical to the green 6G vision. The telecommunications industry can significantly reduce its reliance on fossil fuels and its overall carbon footprint by powering data centers and network operations with solar, wind, and other renewable energies.